OpenStudy (anonymous):

Find the slope of the tangent to the curve y^3x + y^2x^2 = 6 at the point (2,1) I know I need to use derivative, but I've never seen one arranged like this before. Help?

OpenStudy (amistre64):

implicit ....

OpenStudy (amistre64):

what is the derivative of y?

OpenStudy (anonymous):

dy/dx

OpenStudy (amistre64):

yeah, but a simpler notation is y'

OpenStudy (amistre64):

what is the derivative of y^2 ?

OpenStudy (anonymous):

2y?

OpenStudy (amistre64):

and for the record, dy/dx is not correct the derivative of y 'with respect to x' is dy/dx otherwise its just open ended. the derivative of y is just y' now, you are stuck in a certain mindset that the teaching method puts you in for comfort. the derivative of y^2 is a power rule, and a chain rule ... since y is a function of something. (y^2)' = 2y y'

OpenStudy (amistre64):

that is the derivative of: y^2 = 3x ? since the 'with respect to' is not given, make it open ended 2y y' = 3x' when we know its wrt.x, x' = dx/dx = 1 when we know its wrt.y, y' = dy/dy = 1 does this make sense?

OpenStudy (amistre64):

edit: **what is the derivative of ....

OpenStudy (anonymous):

yes that makes sense

OpenStudy (amistre64):

then what is the derivative of y^3 x + y^2 x^2 = 6 [y^3 x]' + [y^2 x^2]' = [6]' (y^3' x + y^3 x') + (y^2' x^2 + y^2 x^2') = 0 (3y^2 y' x + y^3 x') + (2y y' x^2 + y^2 2x x') = 0 solve for y', and we can assume that x' = 1

OpenStudy (anonymous):

Alright thank you so much! It's been so long since we covered this material I forgot how to even begin!

OpenStudy (amistre64):

youre welcome :) y' should now just be a matter of simple algebra, and you know the values of x,y to determine y'

OpenStudy (amistre64):

it may be good to plug in x,y and then solve for y' since the rest will be numbers itll be simpler to pull it out

OpenStudy (amistre64):

3 y' 2 + 1 + 2 y' 2^2 + 2(2) = 0 6y'+ 1 + 8y' + 4 = 0